# Aggregation Models with Optimal Weights for Distributed Gaussian Processes

### v1

@@ 3.3.1 Posterior distribution
where $\mathbf{u}_{\underline{l}}$ is the vector of inducing variables corresponding to the inducing inputs $\Omega_{\underline{l}}$ defined in eq. 18, $\mathbf{K}_{(\cdot)\mathbf{u}_{\underline{l}}}=k(\cdot,\Omega_{\underline{l}})=\mathbf{K}_{\mathbf{u}_{\underline{l}}(\cdot)}^{\top}$, $\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}=k(\Omega_{\underline{l}},\Omega_{\underline{l}})$, and $\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}=\big{[}\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{f}}\mathbf{K}_{\mathbf{f}\mathbf{u}_{\underline{l}}}\big{]}$.
@@ 3.3.2 Optimal coefficients
To compute the optimal coefficients $c_{\underline{l}}$ in GP regression, we need to compute the inner product with respect to the reproducing kernel Hilbert space (RKHS) (Aronszajn,, 1950). Let $\mathcal{H}$ be a RKHS endowed with an inner product $\langle\cdot,\cdot\rangle_{\mathcal{H}}$, then the inner product between functions $f(\cdot)=\sum_{i=1}^{n}\alpha_{i}k(\cdot,\mathbf{x}_{i})$ and $g(\cdot)=\sum_{j=1}^{n^{\prime}}\alpha^{\prime}_{j}k(\cdot,\mathbf{x}^{\prime}_{j})$ is given by
$\langle f,g\rangle_{\mathcal{H}}=\sum_{i=1}^{n}\sum_{j=1}^{n^{\prime}}\alpha_{i}\alpha^{\prime}_{j}k(\mathbf{x}_{i},\mathbf{x}_{j}^{\prime}).$ (36)
In GP regression, the optimization problem in eq. 33 becomes
$R(f)=\sum_{i=1}^{n}(f(\mathbf{x}_{i})-y_{i})^{2}+\lambda\|f\|_{\mathcal{H}}^{2},$ (37)
where $\lambda=\sigma_{\epsilon}^{2}$ is the noise varaince. Here the projection into one of the spaces $V_{\underline{l}}$ has the form
@@
where $\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{k}}}=\big{[}\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{k}}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{f}}\mathbf{K}_{\mathbf{f}\mathbf{u}_{\underline{k}}}\big{]}$. The details are outlined in Algorithm 1 and Algorithm 2 in Appendix.
Figure 1: Errors and time for posterior mean of zero mean and Matérn 3/2 GP with OptiCom and CT of dimension $d=2$ and level $\eta=1,2,\ldots,7$ tested on the Griewank function (Griewank,, 1981) over $n=2000$ training points and $n_{t}=1000$ test points. We denote GP with OptiCom by red dots, GP with CT by blue diamonds. Left: Logarithm of $l_{2}$ errors between GP posteriors and ground truth values averaged over 10 different lengthscales $[0.2,0.4,0.6,0.8,1.0,1.2,1.4,1.6,1.8,2.0]$. Right: Logarithm of time taken to compute the posterior mean averaged over $1000$ replications.
@@ 3.3.3 Complexity
From Lemma 6 in (Garcke,, 2013), we know that the number of summed terms in combination technique formula eq. 27 is $b=\mathcal{O}(\eta^{d-1})$, the size of each summed term is of order $m_{b}=\mathcal{O}(2^{(\eta+d-1)})$, and dimension of the sparse grid space $V_{\eta}^{s}$ defined in eq. 28, i.e., the number of grid points, is $|V_{\eta}^{s}|=\mathcal{O}(2^{\eta}\cdot\eta^{d-1})=\mathcal{O}\big{(}h_{\eta}^{-1}\cdot\log(h_{\eta}^{-1})^{d-1}\big{)}$, where $h_{\eta}=2^{-\eta}$ is the mesh size of the sparse grids $\Omega_{\eta}^{d}$. The additional operations compared to GP with CT include computing $\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{k}}}$ when $\underline{l}\neq\underline{k}$ and solving a $b\times b$ linear system to obtain the optimal coefficients, which costs $\mathcal{O}(b^{2}m_{b}^{2}n)$ and $\mathcal{O}(b^{3})$ time respectively. Note that computation of $\{\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{k}}}\}_{\underline{l},\underline{k}\in\mathbb{I}(\eta,d)}$ and $\{\text{chol}(\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}})\}_{\underline{l}\in\mathbb{I}(\eta,d)}$ can be parallelized and reduced to $\mathcal{O}(m_{b}^{2}n)$ time and $\mathcal{O}(m_{b}^{3})$ time respectively, where $\mathbb{I}(\eta,d)=\{\underline{l}:\eta\leq|\underline{l}|_{1}\leq\eta+d-1\}$. Thus the computational complexity of the parallelized Algorithm 2 is $\mathcal{O}(m_{b}^{3}+b^{3}+m_{b}^{2}n)=\mathcal{O}(b^{3}+m_{b}^{2}n)$. Although the computational complexity of the parallelized GP with OptiCom is larger than that of parallelized GP with CT theoretically, which is $\mathcal{O}(m_{b}^{2}n)$, $b$ is very small in most cases, hence $\mathcal{O}(b^{3}+m_{b}^{2}n)\approx\mathcal{O}(m_{b}^{2}n)$. Figure 1 also demonstrates that GP with OptiCom achieves lower errors and smaller variance than GP with CT at nearly the same time cost.
@@ 4 Distributed GP with optimal weights
In this section, we extend the OptiCom to distributed GPs. Following the assumptions Section 2.2.1, we consider the global training dataset $\mathcal{D}=\{\mathbf{X},\mathbf{y}\}$ of size $n$, partitioned into $M$ local datasets $\mathcal{D}_{i}=\{\mathbf{X}_{i},\mathbf{y}_{i}\}$ of size $n_{i}$ for $i=1,\ldots,M$. All the local datasets are assumed to share the same hyperparameters $\bm{\theta}$.
@@ 4.1.1 Optimal weights
where $f(\cdot)\sim\mathcal{GP}(0,k(\cdot,\cdot^{\prime}))$ as defined in Section 2.2.1, $\mathcal{P}_{i}f=f|\mathbf{y}_{i}:=f_{i}$ is the posterior conditioned on the local dataset $\mathcal{D}_{i}=\{\mathbf{X}_{i},\mathbf{y}_{i}\}$, and its distribution takes the form of Equation 12. Similar to Section 3.3, the optimal weights $\{\beta_{i}\}_{i=1}^{M}$ can be obtained by solving the following linear system:
@@
where $\bm{\alpha}_{l}=[\mathbf{M}_{\mathbf{u}\mathbf{u}}^{(l)}]^{-1}\mathbf{K}_{\mathbf{u}\mathbf{f}_{l}}\mathbf{y}_{l}$, $\mathbf{M}_{\mathbf{u}\mathbf{u}}^{(l)}=\big{[}\mathbf{K}_{\mathbf{u}\mathbf{u}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}\mathbf{f}_{l}}\mathbf{K}_{\mathbf{f}_{l}\mathbf{u}}\big{]}$, $\mathbf{K}_{\mathbf{u}\mathbf{u}}=k(\mathbf{Z},\mathbf{Z})$, $\mathbf{K}_{\mathbf{u}\mathbf{f}_{l}}=k(\mathbf{Z},\mathbf{X}_{l})=\mathbf{K}_{\mathbf{f}_{l}\mathbf{u}}^{\top}$, $\mathbf{M}_{\mathbf{u}\mathbf{u}}^{(c)}=\big{[}\mathbf{K}_{\mathbf{u}\mathbf{u}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}\mathbf{f}_{c}}\mathbf{K}_{\mathbf{f}_{c}\mathbf{u}}\big{]}$, $\mathbf{K}_{\mathbf{u}\mathbf{f}_{c}}=k(\mathbf{Z},\mathbf{X}_{c})=\mathbf{K}_{\mathbf{f}_{c}\mathbf{u}}^{\top}$, $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{m}$ are the inducing inputs corresponding to the optimized inducing variables $\mathbf{u}\in\mathbb{R}^{m}$. The details are outlined in Algorithm 3 in Appendix.
@@ 4.1.2 Aggregated prediction
The details are outlined in Algorithm 4 in Appendix.
@@ 4.1.3 Complexity
Table 1: Complexity of the prediction for aggregation models and full GP. $m$ is the number of inducing points, $n$ is the size of the entire training dataset, $n_{i}$ is the size of the local dataset $\mathcal{D}_{i}$ corresponding to the expert $\mathcal{M}_{i}$, $i=1,\ldots.\,M$, $n_{c}$ is the size of the dataset $\mathcal{D}_{c}$ corresponding to the expert $\mathcal{M}_{c}$ for grBCM. $M$ is the number of experts.
Full PoE gPoE BCM rBCM grBCM NPAE opt (ours)
Time SVGP $\mathcal{O}(m^{2}n)$ $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}(n_{i}+n_{c})+M)$ $\mathcal{O}(m^{2}n_{i}M+M^{3})$ $\mathcal{O}(m^{2}n_{i}+M^{3})$
ExactGP $\mathcal{O}(n^{3})$ $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}((n_{i}+n_{c})^{3}+M)$ $\mathcal{O}(n_{i}^{3}M+M^{3})$ $\mathcal{O}(n_{i}^{3}+M^{3})$
@@ 4.2 Distributed exact GP
$\mu_{\mathcal{A}}(\cdot)=\sum_{i=1}^{M}\beta_{i}\mathbf{K}_{(\cdot)\mathbf{f}_{i}}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{y}_{i},$ (45a)
$\sigma_{\mathcal{A}}^{2}(\cdot)=\sum_{i=1}^{M}\beta_{i}^{2}\mathbf{K}_{(\cdot)\mathbf{f}_{i}}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{K}_{\mathbf{f}_{i}(\cdot)},$ (45b)

### v2

@@ 3.3.1 Posterior distribution
where $\mathbf{u}_{\underline{l}}$ is the vector of inducing variables corresponding to the inducing inputs $\Omega_{\underline{l}}$ defined in eq. 7, $\mathbf{K}_{(\cdot)\mathbf{u}_{\underline{l}}}=k(\cdot,\Omega_{\underline{l}})=\mathbf{K}_{\mathbf{u}_{\underline{l}}(\cdot)}^{\top}$, $\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}=k(\Omega_{\underline{l}},\Omega_{\underline{l}})$, and $\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}=\big[\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{f}}\mathbf{K}_{\mathbf{f}\mathbf{u}_{\underline{l}}}\big]$. To ensure that the predictive covariance remains consistent with that of the full GP, we use the CT coefficients rather than those from OptiCom in the covariance computation.
@@ 3.3.2 Optimal coefficients
To compute the optimal coefficients $c_{\underline{l}}$ in GP regression, we need to compute the inner product with respect to the reproducing kernel Hilbert space (RKHS) (Aronszajn, 1950). Let $\mathcal{H}$ be a RKHS endowed with an inner product $\langle\cdot,\cdot\rangle_{\mathcal{H}}$, then the inner product between functions $f(\cdot)=\sum_{i=1}^{n}\alpha_{i}k(\cdot,\mathbf{x}_{i})$ and $g(\cdot)=\sum_{j=1}^{n^{\prime}}\alpha^{\prime}_{j}k(\cdot,\mathbf{x}^{\prime}_{j})$ is given by $\langle f,g\rangle_{\mathcal{H}}=\sum_{i=1}^{n}\sum_{j=1}^{n^{\prime}}\alpha_{i}\alpha^{\prime}_{j}k(\mathbf{x}_{i},\mathbf{x}_{j}^{\prime})$. In GP regression, the optimization problem in eq. 16 becomes $R(f)=\sum_{i=1}^{n}(f(\mathbf{x}_{i})-y_{i})^{2}+\lambda\|f\|_{\mathcal{H}}^{2}$, where $\lambda=\sigma_{\epsilon}^{2}$ is the noise variance. Here the projection into one of the spaces $V_{\underline{l}}$ has the form
@@
where $\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{k}}}=\big[\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{k}}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{f}}\mathbf{K}_{\mathbf{f}\mathbf{u}_{\underline{k}}}\big]$. The details are outlined in Algorithm 1 and Algorithm 2.
Figure 1: Errors and time for posterior mean of zero mean and Matérn 3/2 GP with OptiCom and CT of dimension $d=2$ and level $\eta=1,2,\ldots,7$ tested on the Griewank function (Griewank, 1981) over $n=2000$ training points and $n_{t}=1000$ test points. We denote GP with OptiCom by circles and GP with CT by diamonds. Left: Logarithm of $l_{2}$ errors between GP posteriors and ground truth values averaged over 10 different lengthscales $[0.2,0.4,0.6,0.8,1.0,1.2,1.4,1.6,1.8,2.0]$. Right: Logarithm of time taken to compute the posterior mean averaged over 10001000 replications.
@@ 3.3.3 Complexity
From Lemma 6 in (Garcke, 2013), we know that the number of summed terms in combination technique formula eq. 11 is $b=\mathcal{O}(\eta^{d-1})$, the size of each summed term is of order $m_{b}=\mathcal{O}(2^{(\eta+d-1)})$, and dimension of the sparse grid space $V_{\eta}^{s}$ defined in eq. 11, i.e., the number of grid points, is $|V_{\eta}^{s}|=\mathcal{O}(2^{\eta}\cdot\eta^{d-1})=\mathcal{O}\big(h_{\eta}^{-1}\cdot\log(h_{\eta}^{-1})^{d-1}\big)$, where $h_{\eta}=2^{-\eta}$ is the mesh size of the sparse grids $\Omega_{\eta}^{d}$. The additional operations compared to GP with CT include computing $\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{k}}}$ when $\underline{l}\neq\underline{k}$ and solving a $b\times b$ linear system to obtain the optimal coefficients, which costs $\mathcal{O}(b^{2}m_{b}^{2}n)$ and $\mathcal{O}(b^{3})$ time respectively. Note that computation of $\{\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{k}}}\}_{\underline{l},\underline{k}\in\mathbb{I}(\eta,d)}$ and $\{\text{chol}(\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}})\}_{\underline{l}\in\mathbb{I}(\eta,d)}$ can be parallelized and reduced to $\mathcal{O}(m_{b}^{2}n)$ time and $\mathcal{O}(m_{b}^{3})$ time respectively, where $\mathbb{I}(\eta,d)=\{\underline{l}:\eta\leq|\underline{l}|_{1}\leq\eta+d-1\}$. Thus the computational complexity of the parallelized Algorithm 2 is $\mathcal{O}(m_{b}^{3}+b^{3}+m_{b}^{2}n)=\mathcal{O}(b^{3}+m_{b}^{2}n)$. Although the computational complexity of the parallelized GP with OptiCom is larger than that of parallelized GP with CT theoretically, which is $\mathcal{O}(m_{b}^{2}n)$, $b$ is very small in most cases, hence $\mathcal{O}(b^{3}+m_{b}^{2}n)\approx\mathcal{O}(m_{b}^{2}n)$. Figure 1 also demonstrates that GP with OptiCom achieves lower errors and smaller variance than GP with CT at nearly the same time cost.
@@ 4 Distributed GP with optimal weights
In our previous formulation of GP with OptiCom, the inducing points were placed on fixed sparse grids rather than being learned through variational inference. While this design simplifies computation, it can limit performance when the sparse grid resolution is low, as the fixed locations may not adequately capture complex input structures. To address this limitation, we extend OptiCom to distributed GPs, which allows for greater flexibility and improved accuracy while maintaining computational efficiency and does not rely on sparse grid structures. Following the assumptions Section 2.2.1, we consider the global training dataset $\mathcal{D}=\{\mathbf{X},\mathbf{y}\}$ of size $n$, partitioned into $M$ local datasets $\mathcal{D}_{i}=\{\mathbf{X}_{i},\mathbf{y}_{i}\}$ of size $n_{i}$ for $i=1,\ldots,M$. All the local datasets are assumed to share the same hyperparameters $\bm{\theta}$.
@@ 4.1.1 Optimal weights
where $f(\cdot)\sim\mathcal{GP}(0,k(\cdot,\cdot))$ as defined in Section 2.2.1, $\mathcal{P}_{i}f=f|\mathbf{y}_{i}:=f_{i}$ is the posterior conditioned on the local dataset $\mathcal{D}_{i}=\{\mathbf{X}_{i},\mathbf{y}_{i}\}$, and its distribution takes the form of eq. 2. Similar to Section 3.3, the optimal weights $\{\beta_{i}\}_{i=1}^{M}$ can be obtained by solving the following linear system:
@@
where $\bm{\alpha}_{l}=[\mathbf{M}_{\mathbf{u}\mathbf{u}}^{(l)}]^{-1}\mathbf{K}_{\mathbf{u}\mathbf{f}_{l}}\mathbf{y}_{l}$, $\mathbf{M}_{\mathbf{u}\mathbf{u}}^{(l)}=\big[\mathbf{K}_{\mathbf{u}\mathbf{u}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}\mathbf{f}_{l}}\mathbf{K}_{\mathbf{f}_{l}\mathbf{u}}\big]$, $\mathbf{K}_{\mathbf{u}\mathbf{u}}=k(\mathbf{Z},\mathbf{Z})$, $\mathbf{K}_{\mathbf{u}\mathbf{f}_{l}}=k(\mathbf{Z},\mathbf{X}_{l})=\mathbf{K}_{\mathbf{f}_{l}\mathbf{u}}^{\top}$, $\mathbf{M}_{\mathbf{u}\mathbf{u}}^{(c)}=\big[\mathbf{K}_{\mathbf{u}\mathbf{u}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}\mathbf{f}_{c}}\mathbf{K}_{\mathbf{f}_{c}\mathbf{u}}\big]$, $\mathbf{K}_{\mathbf{u}\mathbf{f}_{c}}=k(\mathbf{Z},\mathbf{X}_{c})=\mathbf{K}_{\mathbf{f}_{c}\mathbf{u}}^{\top}$, $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{m}$ are the inducing inputs corresponding to the optimized inducing variables $\mathbf{u}\in\mathbb{R}^{m}$. The details are outlined in Algorithm 3.
In addition, the optimal weights computed in eq. 22 inherently capture the correlations between local experts. This mechanism prevents redundant information from being overemphasized when experts are highly correlated, thus leading to a more balanced aggregation. As a result, the aggregated prediction becomes more consistent with that of a full GP. This integration of expert correlations is key to enhancing both the performance and stability of distributed GP.
@@ 4.1.2 Aggregated prediction
The details are outlined in Algorithm 4.
@@ 4.1.3 Complexity
Table 1: Complexity of the prediction for aggregation models. $m$ is the number of inducing points, $n$ is the size of the entire training dataset, $n_{i}$ is the size of the local dataset $\mathcal{D}_{i}$ corresponding to the expert $\mathcal{M}_{i}$, $i=1,\ldots.\,M$, $n_{c}$ is the size of the dataset $\mathcal{D}_{c}$ corresponding to the expert $\mathcal{M}_{c}$ for grBCM. $M$ is the number of experts.
PoE gPoE BCM rBCM grBCM NPAE opt (ours)
Time SVGP $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}(n_{i}+n_{c})+M)$ $\mathcal{O}(m^{2}n_{i}M+M^{3})$ $\mathcal{O}(m^{2}n_{i}+M^{3})$
ExactGP $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}((n_{i}+n_{c})^{3}+M)$ $\mathcal{O}(n_{i}^{3}M+M^{3})$ $\mathcal{O}(n_{i}^{3}+M^{3})$
@@ 4.2 Distributed exact GP
$\mu_{\mathcal{A}}(\cdot)=\sum_{i=1}^{M}\beta_{i}\mathbf{K}_{(\cdot)\mathbf{f}_{i}}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{y}_{i},\quad\sigma_{\mathcal{A}}^{2}(\cdot)=\sum_{i=1}^{M}\beta_{i}^{2}\mathbf{K}_{(\cdot)\mathbf{f}_{i}}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{K}_{\mathbf{f}_{i}(\cdot)}.$ (26)